OpenStudy (anonymous):

Which of these sortition methods can be used to select a person for a certain task from among 24 candidates? Throw 2 dice. Assign a combination (for example 5-6) to each candidate. Flip a coin 5 times. Assign a sequence (for example H-T-H-H-T) to each candidate. Roll a dice 1 time and flip a coin 2 times. Assign a sequence (for example 4-H- T) to each candidate. Flip a coin and roll 2 dice. Assign a sequence (for example H-3-4) to each candidate.

OpenStudy (anonymous):

@unicorn_princess this

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

you are really being asked which experiment has 24 elements in the sample space

OpenStudy (anonymous):

2 dice has \(6\times 6=36\) elements in the sample space, so not that one

OpenStudy (anonymous):

fiipping a coin five times has \(2^5=64\) elements in the sample space, so not that one either

OpenStudy (anonymous):

actually \(2^5=32\) not \(64\) but it is still not 24

OpenStudy (anonymous):

so its d